OpenStudy (anonymous):

What is the area in square feet of a rectangle with a length of 10.6 yards, and a width of 5.8 yards? A. 61.48 ft2 B. 737.76 ft2 C. 184.44 ft2 D. 553.32 ft2

OpenStudy (anonymous):

@J-bird @amistre64 @ganeshie8 @blake57roger @Wolfboy @ShadowLegendX @mathmath333

ganeshie8 (ganeshie8):

Area of rectangle = length x width

ganeshie8 (ganeshie8):

just multiply both the given quantities

ganeshie8 (ganeshie8):

Oh wait, you need to change yards to feet before multiplying

OpenStudy (anonymous):

So 10.6 x 5.8

ganeshie8 (ganeshie8):

that would give you answer in yard^2 but you want answer in ft^2 so convert the given length first

ganeshie8 (ganeshie8):

1 yard = 3 feet

ganeshie8 (ganeshie8):

10.6 yards = 10.6*3 feet = ? 5.8 yards = 5.8*3 feet = ?

OpenStudy (anonymous):

Is it BB

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

nope, try again

ganeshie8 (ganeshie8):

length = 10.6*3 feet width = 5.8*3 feet so area = length x width = 10.6*3 x 5.8*3 = ?

OpenStudy (anonymous):

I got B the first time but now I am getting A

OpenStudy (anonymous):

Am I right or wrong

ganeshie8 (ganeshie8):

use calculator to multiply okay